OpenStudy (anonymous):

Solve for x: 4 over 3x + 2 over 3 = 6

OpenStudy (johnweldon1993):

\[\frac{ 4 }{ 3x } + \frac{ 2 }{ 3 } = 6\] Like that?

OpenStudy (johnweldon1993):

Well if it is like that...you need to first find the common denominator ...which in this case would be 3x....so multiply that second fraction by x/x \[\frac{ 4 }{ 3x } + \frac{ 2x }{ 3x } = 6\] Now put over the common denominator \[\frac{ 4 + 2x }{ 3x }= 6\] And now multiply both sides by 3x to eliminate that fraction \[4 + 2x = 18x\] Now subtract 2x from both sides \[4 = 16x\] And finally divide both sides by 16 to solve for x \[x = \frac{ 4 }{ 16 }\] Simplify \[x = ?\]